# Power-law banded random matrix ensemble as a model for quantum many-body Hamiltonians

### v2

@@ III Basis state labeling
We aim to interpret the PLMRM ensemble as Hamiltonians of quantum many-body systems. This can be accomplished by associating each basis vector with a (different) many-body configuration. We will focus on spin-$\frac{1}{2}$ chains in this work. For simplicity, we restrict to systems without any conservation laws. The Hamiltonian of an $L$-site system is then modeled as a random matrix of dimension $N=2^{L}$, and the many-body configurations (basis states) are strings of $L$ bits. Assigning many-body configurations to basis indices of the random matrix can be done in various ways. We propose and examine three different labeling schemes, which we refer to as ‘random’, binary’, and ‘Gray code’. The binary labeling scheme has been considered before in Refs. [67, 68]. We devise a way to quantify the quality of a labeling scheme, and find the Gray code scheme to be of higher quality than the binary one. We find the (ensemble-averaged) Hamiltonians for the binary and Gray code labeling schemes to be non-homogeneous, meaning that the physical properties are dependent on the site index. We next discuss how to correct for this through what we will refer to as ‘site randomization’.
@@ III.1 Labeling schemes
We represent many-body configurations $c_{i}$ (for example, $c_{i}=001111$) with $i$ ranging over integers from $1$ to $N$ by $L$ binary digits, where the $j$-th digit is zero (one) if the $j$-th spin is in the down (up) state. In power-law banded random matrices, the matrix elements between configurations with neighboring labels have larger magnitude on average. Since physical Hamiltonians contain few-body operators, we thus want neighboring configurations to differ from each other by as few spin flips as possible.
In what follows, we quantify the quality of a given labeling scheme by the ‘badness’
@@
Here, $d_{2}(c_{i},c_{j})$ represents the Hamming distance between the many-body configurations $c_{i}$ and $c_{j}$, i.e., the the minimum number of spin flips connecting the two configurations. Here the subscript $2$ indicates that the numbers are represented in binary digits. The badness penalizes for successive many-body configurations $c_{i}$ and $c_{i+1}$ to be separated by a large number of spin flips.
Figure 2: The badness (7) as a function of the system size $L$ for each of the labeling methods. For the random labeling scheme, the sample-to-sample variance is smaller than the marker size.
@@
A lower badness can be achieved by adapting a labeling scheme that prefers the configurations $c_{i}$ and $c_{i+1}$ to be separated by a smaller number of spin flips. This can be accomplished by taking $c_{i}$ to be the binary representation of $i-1$. We refer to this as the ‘binary’ labeling scheme; this was used in Ref. [67, 68]. For $L=3$, as an example, the labels are
$c_{1}=000\qquad$ $c_{2}=001\qquad$ $c_{3}=010\qquad$ $c_{4}=011$
$c_{5}=100\qquad$ $c_{6}=101\qquad$ $c_{7}=110\qquad$ $c_{8}=111.$
@@
The Gray code provides an ordering of the binary numbers such that successive numbers differ by only a single element [104, 105]. We refer to the corresponding labeling scheme as the ‘Gray code’ labeling scheme, which is characterized by the lowest possible badness $B(L)=1$. The configuration $c_{n}$ in the binary labeling scheme can be converted to the corresponding configuration in the Gray code labeling scheme by taking all up to the first nonzero digit identical. The subsequent digits are equal to $1$ if the corresponding and previous digit in the binary representation are different (i.e., 00 and $1$), and 00 otherwise. Taking $L=3$ again as an example, one finds
$c_{1}=000\qquad$ $c_{2}=001\qquad$ $c_{3}=011\qquad$ $c_{4}=010$
$c_{5}=110\qquad$ $c_{6}=111\qquad$ $c_{7}=101\qquad$ $c_{8}=100.$
We note that there is no unique labeling scheme with the lowest possible badness. (E.g., the badness does not change if the ordering of the digits is changed.) Fig. 2 shows the badness as a function of the system size for the random, binary, and Gray code labeling methods. Clearly, the asymptotic scalings are well-approximated already at small system sizes.
@@ III.2 Site randomization
Figure 3: The ensemble-averaged eigenstate entanglement entropy for mid-spectrum eigenstates with $\alpha=1$, computed using the binary labeling scheme (no site randomization). The spin chain is decomposed in subsystems $A$ and $B$ consisting of the first $L_{A}$ and last $L_{B}=L-L_{A}$ sites, respectively. The inset illustrates the decomposition for $L_{A}=4$ and $L_{B}=6$. The dotted lines give the Page values (8) for $L=6$, $L=10$, and $L=14$.
For $L_{A}/L\approx 1$, the eigenstate entanglement entropy is close to the Page value (dotted line). The entropy is much smaller than the Page value for $L_{A}/L\approx 0$. In order to understand this asymmetry, we start by noting that eigenstates that are not maximally ergodic are not uniformly spread out over all Hilbert space [9]. Such eigenstates tend to have the same binary digits near the left edge (covering subsystem $A$), indicating that they are mainly supported in a specific ‘region’ of the Hilbert space. On the contrary, a change in the digits near the right edge (covering subsystem $B$) changes the resulting basis state number only slightly, such that a basis state in the same region results. Hence, there is no tendency for these digits to be the same, leading to non-symmetric curves when plotting the eigenstate entanglement entropy as a function of $L_{A}/L$. This asymmetry is also seen for the Gray code labeling scheme, and the same argument can be used to explain the effect.
In what follows, every time we use binary labeling or Gray code labeling, we randomize the $L$ real-space site indices, corresponding to the binary digits of the many-body configurations, in order to avoid the (ensemble-averaged) Hamiltonians being non-homogeneous. We refer to this procedure as ‘site randomization’. For the ensemble-averaged results shown below, both the Hamiltonian matrix and the site ordering have been sampled independently for each realization.
Figure 4: The difference between the ensemble-averaged eigenstate entanglement entropy and the Page value as a function of the eigenstate index $n$ (starting from $n=1$ and ordered by increasing energy), scaled by the Hilbert space dimension $N^{L}$. Left and right panels use random and Gray code labeling schemes respectively. The three rows correspond to the three phases ($\alpha=0.25$, $\alpha=0.75$, $\alpha=1.25$). The arrows indicate the direction of flow for the bulk eigenstates with increasing system size. Note the very different scales on the vertical axes for the different values of $\alpha$.
@@ IV Eigenstate Entanglement: Overall Features
We now analyze the many-body interpretations of the PLBRM ensemble. Our focus is on the ensemble-averaged eigenstate entanglement entropy for a bipartition of the system in left and right halves of equal size ($L_{A}=L_{B}=L/2$). Fig. 4 illustrates its dependence on the power-law exponent $\alpha$, the basis state labeling scheme, the system size, and the location on the energy spectrum. The left panels are ensemble-averaged versions of the single-realization scatter plots we presented in Fig. 1. The eigenstate index $n$ starts from $n=1$ and is ordered by increasing energy.
@@
Thus, although all the curves in Fig. 4 nominally have rainbow shapes, only the weakly ergodic phase $\alpha\in\left(\frac{1}{2},1\right)$ shows a true rainbow in the strong sense that the bulk and edge eigenstates have opposite scalings with system size. In the next two subsections we focus separately on the spectral bulk and edge states.
@@ IV.1 Spectral bulk eigenstates
Figure 5: The ensemble-averaged entanglement entropies of mid-spectrum eigenstates, with indices $n$ ranging from $N/2-10$ to $N/2+10$. Top row shows $S_{\text{ent}}$ and bottom row shows $S_{\text{Page}}-S_{\text{ent}}$. The left and right panels use the random and Gray code labeling scheme respectively. In the top row, the horizontal dashed lines give the Page values for the system sizes under consideration.
@@ IV.2 Spectral edge eigenstates
Fig. 6 shows the same plots as discussed above, but now for the spectral-edge eigenstates. For small $\alpha$, we have volume-law ($S_{\text{ent}}\sim L$) scaling as we did for the bulk eigenstates, however, this behavior persists only up to $\alpha\approx\frac{1}{2}$. The crossing separating regimes with volume-law and area-law scaling of $S_{\text{ent}}$ is now located at $\alpha\approx\frac{1}{2}$. (In contrast to $\alpha\approx 1$ seen in Fig. 5 for the bulk eigenstates.) We will study this crossing quantitatively as a function of system size in Sec. V.
@@ V Finite-size dependence
In Figures 5 and 6, we saw crossings near $\alpha\approx 1$ and near $\alpha\approx\frac{1}{2}$ for the bulk and edge eigenstates, respectively. In this section, we present a finite size scaling analysis to determine the value $\alpha_{0}$ of the power-law exponent $\alpha$ at which the scaling of the entanglement entropy changes from volume-law to area-law in the thermodynamic limit. Based on the theory reviewed in Sec. II, and based on the data in Figs. 5 and 6, we expect $\alpha_{0}=1$ for the bulk and $\alpha_{0}=\frac{1}{2}$ for the spectral-edge eigenstates.
@@
Figure 7: The values of $\alpha$ for which $S_{\text{Page}}-S_{\text{ent}}=\gamma$, as a function of $1/L^{2}$ (after ensemble-averaging). The squares, circles, and triangles denote data for the random, binary, and Gray code labeling scheme, respectively. The spectral bulk (left column) and edge (right column) eigenstates are the same as the ones considered in Figs. 5 and 6, respectively. The solid lines give least-square fits of the form (10). The lower panels show the extrapolated value $\alpha_{0}$ for $L\to\infty$ as a function of $\gamma$. The shaded regions in the lower panels indicate the $95\%$ confidence interval of the fit (10).
Fig. 7 shows the values of $\alpha$ at which condition (9) holds as a function of $1/L^{2}$ for both the spectral bulk (left panels) and edge (right panels) eigenstates for $\gamma=0.05$ (top panels) and $\gamma=0.5$ (middle panels). The connecting lines are least-square fits of the second-order polynomial
$\alpha=\alpha_{0}+\alpha_{1}(1/L^{2})+\alpha_{2}(1/L^{2})^{2}$ (10)
with (least-squares) fitting parameters $\alpha_{0}$, $\alpha_{1}$, and $\alpha_{2}$. We observe that these fitted lines connect all data points reasonably well, although the fitting quality is better for the bulk than for the edge. We focus on generic values of $\gamma\in[0.05,0.5]$ in the bottom panels, which show the estimated transition point $\alpha_{0}$ as a function of $\gamma$ together with their uncertainty ($95\%$ confidence interval of the fit (10)). The estimates of $\alpha_{0}$ for the binary and Gray code labeling schemes are near their expected values over the full range. We observe a significant deviation for the random labeling scheme. This means that the entanglement properties of the PLBRM ensemble interpreted as Hamiltonians of quantum many-body systems depend on the choice of the labeling scheme. For small values of $\gamma$, the condition is met for values of $\alpha$ below the crossing, leading to a flow towards larger values with increasing system size. The opposite occurs for large values of $\gamma$.
Figure 8: The difference between the ensemble-averaged eigenstate entanglement entropy and the Page value for the Gray code labeling scheme as a function of the eigenstate index $n$, scaled by $N^{\alpha}$. For clarity, markers are shown only when the horizontal distance between data points is non-negligible.
@@ VI Boundary between bulk and edge
A natural follow-up question is how the transitions at $\alpha\approx\frac{1}{2}$ for the edge and $\alpha\approx 1$ for the bulk eigenstates are related to each other. This relates to the question of how the spectrum is partitioned between bulk and edge parts, at various $\alpha$ in the weakly ergodic phase. We identify two boundaries between different entanglement behaviors in the PLBRM spectrum. In this section, we will restrict to the Gray code labeling scheme.
@@
In the weakly ergodic phase $\frac{1}{2}<\alpha<1$, the bulk eigenstates show volume law entanglement, while the edge eigenstates show area law entanglement. The bulk eigenstates form a measure one fraction of all eigenstates, while the number of edge eigenstates is subextensive. Therefore, we conjecture that $N^{\delta}$ eigenstates at the edges show non-bulk behavior, where $\delta$ is a nontrivial exponent. Depending on the types of entanglement behaviors we want to differentiate, we identify below two separate exponents, which we call $\delta_{1}$ and $\delta_{2}$.
@@
Normalizing the eigenstate index $n$ by $N^{\delta_{1}}$, we looked for a finite-size crossing of curves obtained by plotting $S_{\text{Page}}-S_{\text{ent}}$ as a function of $n/N^{\delta_{1}}$. We find surprisingly good results for $\delta_{1}=\alpha$ (Fig. 8). For any fixed $\alpha$ within the weakly ergodic phase, $\frac{1}{2}<\alpha<1$, we see that the curves for the different system sizes cross at fixed $n/N^{\alpha}\approx 0.2$. The difference $S_{\text{Page}}-S_{\text{ent}}$ increases with system size below this energy and decreases with system size above this energy. The nontrivial exponent $\delta_{1}$ which separates the two behaviors is thus (at least approximately) given by $\delta_{1}=\alpha$. At present we are not aware of an analytic derivation of this result.
@@
Fig. 9 elaborates further on this by showing the ensemble-averaged eigenstate entanglement entropy for the $0.2N^{\alpha}$ eigenstates with the lowest and highest energies as a function of $\alpha$ for several system sizes. The quantity $S_{\text{Page}}-S_{\text{ent}}$ (left panel) shows a crossing at $\alpha\approx 0.5$, while $S_{\text{ent}}/S_{\text{Page}}$ (right panel) shows a crossing at $\alpha\approx 1$. In the weakly ergodic phase $\frac{1}{2}<\alpha<1$, the deviation $S_{\text{Page}}-S_{\text{ent}}$ increases with system size, but the ratio $S_{\text{ent}}/S_{\text{Page}}$ is also non-decreasing. This implies that, in the weakly ergodic phase, the entanglement entropy of the $\approx 0.2N^{\alpha}$ eigenstates at the spectral edges have non-vanishing deviation from $S_{\text{Page}}$, but the majority of these states still have volume-law scaling.
@@
Because $\delta_{1}=\alpha$, when we approach the fully ergodic phase, $\alpha\to\frac{1}{2}+0^{+}$, the number of eigenstates with non-vanishing deviation is $\sim N^{1/2}$. This is a diverging number in the thermodynamic limit. In the fully ergodic phase, $\alpha=\frac{1}{2}+0^{-}$, this number vanishes as all eigenstates have $S_{\text{ent}}\to S_{\text{Page}}$ scaling in this phase. We thus have a discontinuous jump in the number of eigenstates with non-vanishing deviation as we cross the phase boundary at $\alpha=\frac{1}{2}$.
@@
Figure 10: The collapse of the $L=12$ and $L=14$ curves for $\alpha=0.65$ with $\delta_{2}=0.40$ (left) and $\alpha=0.85$ with $\delta_{2}=0.65$ (right) when plotting the ensemble-averaged eigenstate entanglement entropy of the $n$-th eigenstate as a function of $n/N^{\delta_{2}}$.
Figure 11: The scaling of the entanglement entropy of the $N^{\delta}$-th eigenstate as a function of $\alpha$. The coefficient $\delta_{1}$ separates two types of volume-law eigenstates, whereas $\delta_{2}$ separates volume-law eigenstates from area-law eigenstates. The estimates shown for $\delta_{2}$ are lower bounds. The inset shows the estimates of $\delta_{2}$ for collapses between the data for system sizes $L$ and $L+2$.
We uncover a second boundary between edge and bulk behaviors by plotting $S_{\text{ent}}$ directly (instead of plotting $S_{\text{Page}}-S_{\text{ent}}$) against the scaled eigenstate index $n/N^{\delta_{2}}$. If the $S_{\text{ent}}$ versus $n/N^{\delta_{2}}$ curves for different sizes coincide for small $n/N^{\delta_{2}}$ up to a finite value, this indicates an area law ($S_{\text{ent}}\sim L^{0}$) for the lowest $\sim N^{\delta_{2}}$ eigenstates. Thus, we look for a collapse of the $S_{\text{ent}}$ versus $n/N^{\delta_{2}}$ curves for some value of $\delta_{2}$. Fig. 10 shows examples of such collapses. Due to finite-size effects, we can only infer a lower bound for $\delta_{2}$ from our data, as the inferred value of $\delta_{2}$ still increases with increasing system size $L$, for the system sizes available to us. This $L$-dependence is shown in the inset to Fig. 11. The main panel, in turn, shows the lower bound estimate for $\delta_{2}$, obtained from the two largest system sizes at our disposal, $(L,L+2)=(12,14)$.
In order to synthesize these results, in Fig. 11 we show the $\alpha$-dependence of both $\delta_{1}$ and $\delta_{2}$. About $\sim N^{\delta_{2}}$ eigenstates with lowest and highest eigenenergies are area-law states. The number of area-law eigenstates appears to grow from $\mathcal{O}(1)$ at $\alpha=0.5$ to an extensive number at $\alpha=1$. We expect that at $\alpha>1$ most states show area-law entanglement, meaning $\delta_{2}=1$, although from the currently available data we can not explicitly confirm this expectation. There is an intermediate set of ${\sim}N^{\delta_{1}}$ eigenstates which have volume law scaling but a non-vanishing deviation from the Page value. Finally, all the other eigenstates, a measure one fraction, have proper volume law scaling with $S_{\text{ent}}\to S_{\text{Page}}$ in the thermodynamic limit.

### v3

@@ III Basis state labeling
We aim to interpret the PLBRM ensemble as Hamiltonians of quantum many-body systems. This can be accomplished by associating each basis vector with a (different) many-body configuration. We will focus on spin-$\frac{1}{2}$ chains in this work. For simplicity, we restrict to systems without any conservation laws. The Hamiltonian of an $L$-site system is then modeled as a random matrix of dimension $N=2^{L}$, and the many-body configurations (basis states) are strings of $L$ bits. Assigning many-body configurations to basis indices of the random matrix can be done in various ways. We propose and examine three different labeling schemes, which we refer to as ‘random’, binary’, and ‘Gray code’. The binary labeling scheme has been considered before in Refs. [68, 69]. We devise a way to quantify the quality of a labeling scheme, and find the Gray code scheme to be of higher quality than the binary one. We find the (ensemble-averaged) Hamiltonians for the binary and Gray code labeling schemes to be non-homogeneous, meaning that the physical properties are dependent on the site index. We next discuss how to correct for this through what we will refer to as ‘site randomization’.
@@ III.1 Labeling schemes
We represent many-body configurations $c_{i}$ (for example, $c_{i}=001111$) with $i$ ranging over integers from $1$ to $N$ by $L$ binary digits, where the $j$th digit is zero (one) if the $j$th spin is in the down (up) state. In power-law banded random matrices, the matrix elements between configurations with neighboring labels have larger magnitude on average. Since physical Hamiltonians contain few-body operators, we thus want neighboring configurations to differ from each other by as few spin flips as possible.
In what follows, we quantify the quality of a given labeling scheme by the “badness”
@@
Here, $d_{2}(c_{i},c_{j})$ represents the Hamming distance between the many-body configurations $c_{i}$ and $c_{j}$, i.e., the minimum number of spin flips connecting the two configurations. Here the subscript $2$ indicates that the numbers are represented in binary digits. The badness penalizes for successive many-body configurations $c_{i}$ and $c_{i+1}$ to be separated by a large number of spin flips.
Figure 2: Badness (7) as a function of the system size $L$ for each of the labeling methods. For the random labeling scheme, the sample-to-sample variance is smaller than the marker size.
@@
A lower badness can be achieved by adapting a labeling scheme that prefers the configurations $c_{i}$ and $c_{i+1}$ to be separated by a smaller number of spin flips. This can be accomplished by taking $c_{i}$ to be the binary representation of i1i-1. We refer to this as the ‘binary’ labeling scheme; this was used in Ref. [68, 69]. For $L=3$, as an example, the labels are
$c_{1}=000,\qquad$ $c_{2}=001,\qquad$ $c_{3}=010,\qquad$ $c_{4}=011,$
$c_{5}=100,\qquad$ $c_{6}=101,\qquad$ $c_{7}=110,\qquad$ $c_{8}=111.$
@@
The Gray code provides an ordering of the binary numbers such that successive numbers differ by only a single element [105, 106]. We refer to the corresponding labeling scheme as the “Gray code” labeling scheme, which is characterized by the lowest possible badness $B(L)=1$. The configuration $c_{n}$ in the binary labeling scheme can be converted to the corresponding configuration in the Gray code labeling scheme by taking all up to the first nonzero digit identical. The subsequent digits are equal to $1$ if the corresponding and previous digit in the binary representation are different (i.e., 0 and $1$), and 0 otherwise. Taking $L=3$ again as an example, one finds
$c_{1}=000,\qquad$ $c_{2}=001,\qquad$ $c_{3}=011,\qquad$ $c_{4}=010,$
$c_{5}=110,\qquad$ $c_{6}=111,\qquad$ $c_{7}=101,\qquad$ $c_{8}=100.$
We note that there is no unique labeling scheme with the lowest possible badness (e.g., the badness does not change if the ordering of the digits is changed). Figure 2 shows the badness as a function of the system size for the random, binary, and Gray code labeling methods. Clearly, the asymptotic scalings are well-approximated already at small system sizes.
@@ III.2 Site randomization
Figure 3: Ensemble-averaged eigenstate entanglement entropy for mid-spectrum eigenstates with $\alpha=1$, computed using the binary labeling scheme (no site randomization). The spin chain is decomposed in subsystems $A$ and $B$ consisting of the first $L_{A}$ and last $L_{B}=L-L_{A}$ sites, respectively. The inset illustrates the decomposition for $L_{A}=4$ and $L_{B}=6$. The dotted lines give the Page values (8) for $L=6$, $L=10$, and $L=14$.
For $L_{A}/L\approx 1$, the eigenstate entanglement entropy is close to the Page value (dotted line). The entropy is much smaller than the Page value for $L_{A}/L\approx 0$. To understand this asymmetry, we start by noting that eigenstates that are not maximally ergodic are not uniformly spread out over all Hilbert space [9]. Such eigenstates tend to have the same binary digits near the left spatial edge (covering subsystem $A$), indicating that they are mainly supported in a specific ‘region’ of the Hilbert space. On the contrary, a change in the digits near the right spatial edge (covering subsystem $B$) changes the resulting basis state number only slightly, meaning that the state is in the same Hilbert space region before and after the change in the digits. Hence, there is no tendency for these digits to be the same, leading to non-symmetric curves when plotting the eigenstate entanglement entropy as a function of $L_{A}/L$. This asymmetry is also seen for the Gray code labeling scheme, and the same argument can be used to explain the effect. We stress that the discussion here is about the spatial edges, not the spectral ones.
In what follows, every time we use binary labeling or Gray code labeling, we randomize the $L$ real-space site indices, corresponding to the binary digits of the many-body configurations, to avoid the (ensemble-averaged) Hamiltonians being nonhomogeneous. We refer to this procedure as “site randomization”. For the ensemble-averaged results shown below, both the Hamiltonian matrix and the site ordering have been sampled independently for each realization.
Figure 4: Difference between the ensemble-averaged eigenstate entanglement entropy and the Page value as a function of the eigenstate index $n$ (starting from $n=1$ and ordered by increasing energy), scaled by the Hilbert space dimension NN. On the horizontal axes, the ground states are on the left, while the highest excited eigenstates are on the right. Left and right panels use random and Gray code labeling schemes respectively. The three rows correspond to the three phases ($\alpha=0.25$, $\alpha=0.75$, $\alpha=1.25$). The arrows indicate the direction of flow for the bulk eigenstates with increasing system size. Note the very different scales on the vertical axes for the different values of $\alpha$.
@@ IV Eigenstate Entanglement: Overall Features
We now analyze the many-body interpretations of the PLBRM ensemble. Our focus is on the ensemble-averaged eigenstate entanglement entropy for a bipartition of the system in left and right halves of equal size ($L_{A}=L_{B}=L/2$). Fig. 4 illustrates its dependence on the power-law exponent $\alpha$, the basis state labeling scheme, the system size, and the location on the energy spectrum. The left panels are ensemble-averaged versions of the single-realization scatter plots we presented in Fig. 1. The eigenstate index $n$ starts from $n=1$ and is ordered by increasing energy, meaning that the ground states are on the left, while the highest excited eigenstates are on the right of the horizontal axes.
@@
Thus, although all the curves in Fig. 4 nominally have rainbow shapes, only the weakly ergodic phase $\alpha\in\left(\frac{1}{2},1\right)$ shows a true rainbow in the strong sense that the spectral bulk and spectral edge eigenstates have opposite scalings with system size. In the next two subsections we focus separately on the spectral bulk and edge eigenstates.
@@ IV.1 Spectral bulk eigenstates
Figure 5: Ensemble-averaged entanglement entropies of mid-spectrum eigenstates, with indices $n$ ranging from N/210N/2-10 to $N/2+10$. Top row shows $S_{\text{ent}}$ and bottom row shows $S_{\text{Page}}-S_{\text{ent}}$. The left and right panels use the random and Gray code labeling scheme respectively. In the top row, the horizontal dashed lines give the Page values for the system sizes under consideration.
@@ IV.2 Spectral edge eigenstates
Figure 6 shows the same plots as discussed above, but now for the spectral-edge eigenstates. For small $\alpha$, we have volume-law ($S_{\text{ent}}\sim L$) scaling as we did for the bulk eigenstates, however, this behavior persists only up to $\alpha\approx\frac{1}{2}$. The crossing separating regimes with volume-law and area-law scaling of $S_{\text{ent}}$ is now located at $\alpha\approx\frac{1}{2}$. (In contrast to $\alpha\approx 1$ seen in Fig. 5 for the bulk eigenstates.) We will study this crossing quantitatively as a function of system size in Section V.
@@ V Finite-size dependence
In Figures 5 and 6, we saw crossings near $\alpha\approx 1$ and near $\alpha\approx\frac{1}{2}$ for the spectral bulk and spectral edge eigenstates, respectively. In this section, we present a finite size scaling analysis to determine the value $\alpha_{0}$ of the power-law exponent $\alpha$ at which the scaling of the entanglement entropy changes from volume-law to area-law in the thermodynamic limit. Based on the theory reviewed in Sec. II, and based on the data in Figs. 5 and 6, we expect $\alpha_{0}=1$ for the spectral bulk and $\alpha_{0}=\frac{1}{2}$ for the spectral-edge eigenstates.
@@
Figure 7: Values of $\alpha$ for which $S_{\text{Page}}-S_{\text{ent}}=\gamma$, as a function of $1/L^{2}$ (after ensemble-averaging). The squares, circles, and triangles denote data for the random, binary, and Gray code labeling scheme, respectively. The spectral bulk (left column) and edge (right column) eigenstates are the same as the ones considered in Figs. 5 and 6, respectively. The solid lines give least-square fits of the form (10). The lower panels show the extrapolated value $\alpha_{0}$ for $L\to\infty$ as a function of $\gamma$. The shaded regions in the lower panels indicate the $95\%$ confidence interval of the fit (10).
Figure 7 shows the values of $\alpha$ at which condition (9) holds as a function of $1/L^{2}$ for both the spectral bulk (left panels) and edge (right panels) eigenstates for $\gamma=0.05$ (top panels) and $\gamma=0.5$ (middle panels). The connecting lines are least-square fits of the second-order polynomial
$\alpha=\alpha_{0}+\alpha_{1}(1/L^{2})+\alpha_{2}(1/L^{2})^{2},$ (10)
with (least-squares) fitting parameters $\alpha_{0}$, $\alpha_{1}$, and $\alpha_{2}$. We observe that these fitted lines connect all data points reasonably well, although the fitting quality is better for the spectral bulk than for the spectral edge eigenstates. We focus on generic values of $\gamma\in[0.05,0.5]$ in the bottom panels, which show the estimated transition point $\alpha_{0}$ as a function of $\gamma$ together with their uncertainty $95\%$ confidence interval of the fit (10)]. The estimates of $\alpha_{0}$ for the binary and Gray code labeling schemes are near their expected values over the full range. We observe a significant deviation for the random labeling scheme. This means that the entanglement properties of the PLBRM ensemble interpreted as Hamiltonians of quantum many-body systems depend on the choice of the labeling scheme. For small values of $\gamma$, the condition is met for values of $\alpha$ below the crossing, leading to a flow towards larger values with increasing system size. The opposite occurs for large values of $\gamma$.
Figure 8: Difference between the ensemble-averaged eigenstate entanglement entropy and the Page value for the Gray code labeling scheme as a function of the eigenstate index $n$, scaled by $N^{\alpha}$. The insets show the same quantity as a function of $n/N^{\delta}$ for $\delta=\alpha+0.1$. For clarity, markers are shown only when the horizontal distance between data points is non-negligible.
@@ VI Boundary between bulk and edge
A natural follow-up question is how the transitions at $\alpha\approx\frac{1}{2}$ for the spectral edge and $\alpha\approx 1$ for the spectral bulk eigenstates are related to each other. This relates to the question of how the spectrum is partitioned between bulk and edge parts, at various $\alpha$ in the weakly ergodic phase. We identify two boundaries between different entanglement behaviors in the PLBRM spectrum. In this section, we will restrict to the Gray code labeling scheme.
@@
In the weakly ergodic phase $\frac{1}{2}<\alpha<1$, the spectral bulk eigenstates show volume law entanglement, while the spectral edge eigenstates show area law entanglement. The spectral bulk eigenstates form a measure one fraction of all eigenstates, while the number of spectral edge eigenstates is subextensive. Naturally, one expects the number of edge eigenstates (i) to increase with $N$, and (ii) to become a zero fraction in the thermodynamic limit. Although various scaling forms obey these properties, $N^{\delta}$ with $\delta<1$ is arguably the simplest form. Therefore, we hypothesize that $N^{\delta}$ eigenstates at the spectral edges show non-bulk behavior, where $\delta$ is a nontrivial exponent. Depending on the types of entanglement behaviors we want to differentiate, we introduce below two separate exponents, which we call $\delta_{1}$ and $\delta_{2}$.
@@
Normalizing the eigenstate index $n$ by $N^{\delta_{1}}$, we looked for a finite-size crossing of curves obtained by plotting $S_{\text{Page}}-S_{\text{ent}}$ as a function of $n/N^{\delta_{1}}$. We find surprisingly good results for $\delta_{1}=\alpha$ (Fig. 8). For any fixed $\alpha$ within the weakly ergodic phase, $\frac{1}{2}<\alpha<1$, we see that the curves for the different system sizes cross at fixed $n/N^{\alpha}\approx 0.2$. The difference $S_{\text{Page}}-S_{\text{ent}}$ increases with system size below this energy and decreases with system size above this energy. When $n$ is scaled by $N^{\delta}$ with $\delta\neq\alpha$, the curves no longer cross at approximately the same point, as shown in the insets for $\delta=\alpha+0.1$. The nontrivial exponent $\delta_{1}$ which separates the two behaviors is thus (at least approximately) given by $\delta_{1}=\alpha$. At present we are not aware of an analytic derivation of this result.
@@
Fig. 9 elaborates further on this by showing the ensemble-averaged eigenstate entanglement entropy for the $0.2N^{\alpha}$ eigenstates with the lowest and highest energies as a function of $\alpha$ for several system sizes. The quantity $S_{\text{Page}}-S_{\text{ent}}$ (left panel) shows a crossing at $\alpha\approx 0.5$, while $S_{\text{ent}}/S_{\text{Page}}$ (right panel) shows a crossing at $\alpha\approx 1$. In the weakly ergodic phase $\frac{1}{2}<\alpha<1$, the deviation $S_{\text{Page}}-S_{\text{ent}}$ increases with system size, but the ratio $S_{\text{ent}}/S_{\text{Page}}$ is also non-decreasing. This implies that, in the weakly ergodic phase, the entanglement entropy of the $\approx 0.2N^{\alpha}$ eigenstates at the spectral edges have nonvanishing deviation from $S_{\text{Page}}$, but the majority of these states still have volume-law scaling.
@@
Because $\delta_{1}=\alpha$, when we approach the fully ergodic phase, $\alpha\to\frac{1}{2}+0^{+}$, the number of eigenstates with nonvanishing deviation is $\sim N^{1/2}$. This is a diverging number in the thermodynamic limit. In the fully ergodic phase, $\alpha=\frac{1}{2}+0^{-}$, this number vanishes as all eigenstates have $S_{\text{ent}}\to S_{\text{Page}}$ scaling in this phase. We thus have a discontinuous jump in the number of eigenstates with nonvanishing deviation as we cross the critical point at $\alpha=\frac{1}{2}$.
@@
Figure 10: Collapse of the $L=12$ and $L=14$ curves for $\alpha=0.65$ with $\delta_{2}=0.40$ (left) and $\alpha=0.85$ with $\delta_{2}=0.65$ (right) when plotting the ensemble-averaged eigenstate entanglement entropy of the $n$-th eigenstate as a function of $n/N^{\delta_{2}}$.
Figure 11: Scaling of the entanglement entropy of the $N^{\delta}$-th eigenstate as a function of $\alpha$. The coefficient $\delta_{1}$ separates two types of volume-law eigenstates, whereas $\delta_{2}$ separates volume-law eigenstates from area-law eigenstates for the largest studied system size. The estimates shown for $\delta_{2}$ are lower bounds. The inset shows the estimates of $\delta_{2}$ for collapses between the data for system sizes $L$ and $L+2$.
We uncover a second boundary between edge and bulk behaviors by plotting $S_{\text{ent}}$ directly (instead of plotting $S_{\text{Page}}-S_{\text{ent}}$) against the scaled eigenstate index $n/N^{\delta_{2}}$. If the $S_{\text{ent}}$ versus $n/N^{\delta_{2}}$ curves for different sizes coincide for small $n/N^{\delta_{2}}$ up to a finite value, this indicates an area law ($S_{\text{ent}}\sim L^{0}$) for the lowest $\sim N^{\delta_{2}}$ eigenstates. Thus, we look for a collapse of the $S_{\text{ent}}$ versus $n/N^{\delta_{2}}$ curves for some value of $\delta_{2}$. Figure 10 shows examples of such collapses. Due to finite-size effects, we can only infer a lower bound for $\delta_{2}$ from our data, as the inferred value of $\delta_{2}$ still increases with increasing system size $L$, for the system sizes available to us. This $L$-dependence is shown in the inset to Fig. 11. The main panel, in turn, shows the lower bound estimate for $\delta_{2}$, obtained from the two largest system sizes at our disposal, $(L,L+2)=(12,14)$.
To synthesize these results, in Fig. 11 we show the $\alpha$-dependence of both $\delta_{1}$ and $\delta_{2}$. About $\sim N^{\delta_{2}}$ eigenstates with lowest and highest eigenenergies are area-law states. The number of area-law eigenstates appears to grow from $\mathcal{O}(1)$ at $\alpha=0.5$ to an extensive number at $\alpha=1$. We expect that at $\alpha>1$ most eigenstates show area-law entanglement, meaning $\delta_{2}=1$, although from the currently available data we can not explicitly confirm this expectation. There is an intermediate set of ${\sim}N^{\delta_{1}}$ eigenstates which have volume law scaling but a nonvanishing deviation from the Page value. Finally, all the other eigenstates, a measure one fraction, have proper volume law scaling with $S_{\text{ent}}\to S_{\text{Page}}$ in the thermodynamic limit.